Exercise 52: A consumer watchdog organization is concerned about credit card debt. A survey of 10 young
adults with credit card debt of more than $2,000 showed they paid an average of just over
$100 per month against their balances. Listed below is the amounts each young adult paid
last month. 110 116 103 93 99
113 87 101 109 104
Determine the sample standard deviation. [answer is the same using STDEV or STDEV.S]
Exercise 54: The mean of a group of sample observations is 500 & the standard deviation is 50.  
According to Chebyshev's theorem, at least what percentage of the data will lie between 400 & 600?
Answer #1

52)The sample mean is

+ X10 10

The sample standard deviation is

2 = 8.9969 10-1

54)The Chebyshev's inequality is

σ2

Which is equivalent to,

σ2 1P1 σ2

We need to find

P(400 < X 〈 600) = P(400-500 < X-500 < 600-500) P(400 < X 〈 600) = P(-100 < X-500 < 100) P(400 < X < 600)-P-100 < X-μ < 100)

Now applying Chebyshev bound,

P(400 〈 X 〈 600) = PCIX-μ| < 100) 502 P(400 . X 〈 600) > 1-1002 3 P(400 < X < 600) > 4

So at least 75\% of data will lie between 400 and 600.